OpenStudy (anonymous):

Use the distributive property to simplify the expression below: -5x + 20 Divided by -5

OpenStudy (anonymous):

Do each term seperately.\[\frac{ -5x + 20 }{ -5 } = \frac{ -5x }{ -5 } + \frac{ 20 }{ -5 }\]-5x divided by -5 = 1x or x 20 divided by -5 is -4. So the simplified form of the expression is x - 4.

OpenStudy (anonymous):

Okay thats what i thought thank you soo much.

OpenStudy (anonymous):

No problem. (: